# MF-LAL: Drug Compound Generation Using Multi-Fidelity Latent Space Active Learning

### v1

@@ 5 Discussion and Conclusion
Limitations of our approach include a limited set of oracles and a potential lack of synthesizability of the generated compounds, since SAscore is known to be imperfect (Skoraczyński et al., 2023). Future work could include using a reaction-aware generative model that generates more synthesizable molecules (Horwood & Noutahi, 2020).
6 Acknowledgments
This work was supported in part by Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, CDC-RFA-FT-23-0069, NSF Grants #2205093, #2146343, and #2134274. MKG has an equity interest in and is a cofounder and scientific advisor of VeraChem LLC.
@@ Appendix A Model details
A.1 MF-LAL model diagram
Figure 3: Diagram of MF-LAL architecture. An input molecule is encoded into the first latent space $\mathbf{z}_{1}$ using a neural network. The networks $h_{\xi_{1}},\dots,h_{\xi_{K-1}}$ transform points in $z_{1}$ to the higher fidelity latent spaces. Each latent space has an associated decoder, which reconstructs the original molecule, and a GP surrogate model for that fidelity level.
@@ Appendix B Experimental details
All experiments were conducted on a server with 8 RTX 2080 Ti GPUs. For our model and each baseline, we performed a random hyperparameter search with 20 trials using only the first 3 fidelities, and took the set of hyperparameters with the best generated samples at $f_{3}$ following 3 hours of active learning. We excluded ABFE from the hyperparameter search due to computational cost, and just used the same set of hyperparameters for all subsequent experiments using all 4 fidelities.
@@ B.1 Oracles
For all oracles, we estimated the cost (for the baselines that require it) using the average run time over 10 samples with random input compounds. We also computed the ROC-AUC (shown below) of each oracle for the BRD4(2) target to confirm that the higher cost oracles are more accurate. To do this, we generated a dataset of known active and presumed inactive compounds for BRD4(2), and then ran each oracle on all of the compounds. The actives were retrieved from the BRD4(2) target from BindingDB with $K_{i}<10\mu M$, and the inactives were generated using DUD-E (Mysinger et al., 2012). As expected, the ROC-AUC increases with higher fidelity, as well as the computational cost.
Linear regression (cost: 0.1s, ROC-AUC: 0.59)
@@
Ensembled AutoDock4 (cost: 44s, ROC-AUC: 0.70)
@@
Absolute binding free energy (ABFE) (cost: 9:20hrs, ROC-AUC: 0.92)

### v3

@@ 5 Discussion and Conclusion
Limitations of our approach include a limited set of oracles and a potential lack of synthesizability of the generated compounds, since SAscore is known to be imperfect (Skoraczyński et al., 2023). The SVGP technique we use for our surrogate models is known to overestimate the posterior variance far away from the inducing points (Bauer et al., 2016), potentially biasing our method towards out-of-distribution molecules and increasing the unpredictability of the surrogate outputs. We also note that by design, MF-LAL may miss scaffolds at the highest fidelity that have strong binding if the lower fidelity oracles did not score that scaffold well. We consider this a necessary tradeoff to make the search for good compounds at the highest fidelity computationally tractable. Finally, it should also be noted that our experimental setup of evaluating each method after a fixed run time of 1 week may lead to high variance in the results, because each method has not been trained to convergence. While this most accurately reflects real-world usage, it may mean the ABFE values of the top generated compounds are somewhat variable between runs. Future work could include using a reaction-aware generative model that generates more synthesizable molecules (Horwood & Noutahi, 2020).
Impact Statement
Similar to other works that apply machine learning to drug discovery, our work is subject to dual use (Urbina et al., 2022). There is potential for societal benefit, by helping develop new drug compounds to treat disease. However, there is also potential for harm, such as to generate new chemical weapons. Fortunately, the latter places a whole additional set of requirements on compounds (e.g. skin-absorbable or volatile and subject to inhalation), so this problematic direction does not appear to be imminent.
Acknowledgments
This work was supported in part by NSF Grants #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069, the U.S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, DARPA AIE FoundSci and DARPA YFA. MKG has an equity interest in and is a cofounder and scientific advisor of VeraChem LLC.
@@ Appendix A Model details
A.1 Training details
Given a training molecule $x$ and associated oracle output $y$ at fidelity $k$, we minimize
[…]
where the first term is equivalent to maximizing the ELBO and the second is the MLL of the GP. While the loss is only evaluated at fidelity $k$, it is backpropagated through to all lower fidelities.
For the likelihood term of the generation objective, we maximize the likelihood of a point $\mathbf{z}^{(i)}_{k}$ in latent space $k$. This is equivalent to maximizing the probability density function evaluated at that point:
Additionally, we add a diversity term to the generation objective. We aim to minimize the following function:
where $s(A,B)$ is the cosine similarity between vectors $A$ and $B$.
@@ Appendix B Experimental details
All experiments were conducted on a server with 8 RTX 2080 Ti GPUs. For our model and each baseline, we performed a random hyperparameter search with 20 trials using only the first 3 fidelities, and took the set of hyperparameters with the best generated samples at $f_{3}$ following 3 hours of active learning. We excluded ABFE from the hyperparameter search due to computational cost, and just used the same set of hyperparameters for all subsequent experiments using all 4 fidelities. We note that MF-LAL was somewhat sensitive to choice of hyperparameters, especially the KLD/reconstruction ratio for the decoders and the diversity coefficient used during generation. Thus, we advise users of MF-LAL to conduct a hyperparameter search similar to the method we used before applying MF-LAL to a new set of oracles.
@@ B.1 Oracles
For all oracles, we estimated the cost (for the baselines that require it) using the average run time over 10 samples with random input compounds. We also computed the ROC-AUC (shown below) of each oracle for the BRD4(2) and c-MET targets to confirm that the higher cost oracles are more accurate. To do this, we generated a dataset of 32 known active and presumed inactive compounds for BRD4(2) and c-MET, and then ran each oracle on all of the compounds. The actives were retrieved from the BRD4(2)/c-MET target from BindingDB with $K_{i}<10\mu M$, and the inactives were generated using DUD-E (Mysinger et al., 2012). As expected, for both targets, the ROC-AUC increases with higher fidelity, as well as the computational cost. This shows that the higher cost oracles are indeed more accurate, and so the hierarchy of fidelities we use is sensible and applicable to multi-fidelity learning.
Linear regression (cost: 0.1s, ROC-AUC BRD4(2): 0.59, c-MET: 0.68)
@@
Ensembled AutoDock4 (cost: 44s for BRD4(2) and 68s for c-MET, ROC-AUC BRD4(2): 0.80, c-MET: 0.80)
@@
Absolute binding free energy (ABFE) (cost: 9:20hrs, ROC-AUC BRD4(2): 0.92, c-MET: 0.89)
@@ B.2 Baselines
  • RGA (only docking) (Fu et al., 2022). Used the code available at https://github.com/futianfan/reinforced-genetic-algorithm.
@@
  • MF-GP + ZINC250k. We use the Multi-Fidelity Max Value Entropy acquisition function for selecting compounds during active learning (Takeno et al., 2020), and a linear truncated fidelity kernel (Gardner et al., 2018) for the GP surrogate. All ZINC250k (Irwin et al., 2012) compounds are provided as the candidate set.
@@
  • DecompDiff (Guan et al., 2023b). Used the code available at https://github.com/bytedance/DecompDiff.
  • TAGMol (only docking) (Dorna et al., 2024). Used the code available at https://github.com/MoleculeAI/TAGMol. We used the single-objective gradient guidance model for optimizing binding affinity.
B.3 Statistical tests
To compare the mean ABFE score of generated compounds from each method, we used a standard Student’s t-test. We also used a standard t-test for the mean of the top 3 generated compounds from each method. For these tests, we used the entire set of 40 compounds if we had run them for a given method, otherwise we used 15. We only analyzed the differences in the top 3 compounds among the methods that had 40 compounds. For comparing the number of active scaffolds, we used a binomial test to compare the proportion of distinct active compounds among the total generated compounds. For this test, we cannot directly compare the samples of 40 compounds with the samples of 15 compounds, because we expect the number of active scaffolds to plateau as we generate more compounds (meaning a proportion test across samples with different sizes is not valid). Therefore, the p-value we report for number of active scaffolds only compares MF-LAL with the top baseline. However, we found that limiting all methods to 15 generated compounds still showed statistically significant results favoring MF-LAL.
@@ C.1 Oracle outputs during active learning
C.2 Reconstruction accuracy
Figure 5: Reconstruction accuracy during active learning. The y-axis shows the proportion of training set compounds that were successfully reconstructed using the decoder, and the x-axis shows the active learning iteration.
Figure 5 shows the reconstruction accuracy of the decoders during active learning. Reconstruction accuracy is the proportion of compounds that were exactly reconstructed, meaning every SELFIES character is identical to the input. The decoder corresponding to the highest fidelity latent space achieves the overall highest reconstruction accuracy, which is likely because it only has to decode from a very limited compound space. The lowest fidelity decoder has the worst reconstruction accuracy, because it decodes the most varied set of compounds. Nonetheless, the reconstruction accuracies are relatively high across all decoders, meaning MF-LAL successfully learns the mapping from latent space back to molecule at all fidelities.